OpenStudy (anonymous):

Please solve the equation on the interval 0 (less-than or equal-to) theta < 2pi sin(2theta) = -(sqrt3)/2 Thank you

OpenStudy (anonymous):

If you add fish to the equation then you end up with cat x 7 / 3

OpenStudy (anonymous):

5pi/6

OpenStudy (campbell_st):

well using tri ratios \[2\theta = \sin^{-1} (\sqrt{3}/2) = \pi/3\] sin is negative in 3rd quadrant ( pi + theta) and 4th quadrant ( 2pi - theta) so the solutions are \[\theta = 4\pi/3, 5\pi/3\]

OpenStudy (anonymous):

2θ = π/3 --> θ = π/6

OpenStudy (anonymous):

How did I get 2pi/3 and 5pi/3

OpenStudy (campbell_st):

oops should be pi/6.... forgot to have it

OpenStudy (campbell_st):

sin is negative... thats the key............. in the 3rd quadrant and 4th quadrant sin(x) is negative... so find theta then 3rd quadrant is (pi + theta) 4th quadrant is (2pi - theta)

OpenStudy (anonymous):

Oh no...i don't believe it, i typed the equation incorrectly....sqrt3/2 is not negative, it is positive

OpenStudy (anonymous):

im so sorry

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

That's why you have different answer from ours!

OpenStudy (campbell_st):

type into your calculator sin(7pi/6) and (11pi/6) is the answer \[-\sqrt{3}/2\]

OpenStudy (campbell_st):

then its 1st and 2nd quadrants where sin is positive so the angles are pi/6 and pi - pi/6 = 5pi/6

OpenStudy (anonymous):

i got theta=pi/6 and 7pi/6

OpenStudy (anonymous):

@Mike, if you're happy with the solver, click on the blue Best answer, will you :)

OpenStudy (campbell_st):

ummmm good luck

OpenStudy (anonymous):

thanks